2. a) One day, Barry (assuming one man in the economy), the barber collects Tk. 400 for haircuts. Over this day, his equipment depreciates in value by Tk.50. Of the remaining Tk. 350, Barry sends Tk. 30 to the government in sales taxes, takes home Tk. 220 in wages, and retains Tk. 100 in his business to add new equipment in the future. From the Tk. 220 that barry takes home, he pays Tk. 70 in income taxes. Based on this information compute - Gross domestic product (GDP), Net national Product (NNP), National income (NI), Personal Income (PI), Disposable personal income.
1
Expert's answer
2020-04-15T09:18:41-0400

"GDP=Tk.400."

"NNP=GDP\u2212depreciation=400\u221250=Tk.350."

"NI=GDP+foreighincomes=400+0=Tk.400."

"PI=350-30-100=Tk.220."

"DI= PI-Taxes= 220-70= Tk.150."